Suppose a country has a national debt of $5,000 billion, a GDP of $10,000 billion, and a budget deficit of $100 billion.

1. How much will its new national debt be?
2. Compute its debt-GDP ratio.
3. Suppose its GDP grows by 1% in the next year and the budget deficit is again $100 billion. Compute its new level of national debt and its new debt-GDP ratio.
4. Explain your answers for all of these questions (one well composed paragraph for each question).
1
Expert's answer
2014-04-02T08:59:40-0400
National debt = $5,000 billion, GDP = $10,000 billion, budget deficit = $100 billion.
1. National debt = National debt of the previous year + budget deficit = 5,000 + 100 = $5,100 billion, so the national debt increased.
2. Debt-GDP ratio = National debt/GDP*100% = 5,000/10,000*100% = 50% - the debt is big enough but is not critical.
3. Suppose its GDP grows by 1% in the next year and the budget deficit is again $100 billion. Compute its new level of national debt and its new debt-GDP ratio.
New GDP = 10,000*1.01 = $10,100 billion
New level of national debt = 5,100 + 100 = $5,200 billion
New debt-GDP ratio = 5,200/10,100*100% = 51.5%.
4. So, as we can see, both national debt and debt-GDP ratio increased, so the government should try to decrease its budget deficit.